# Advances in the Shannon Capacity of Graphs

### v1

@@ 3 The Shannon capacity of polynomials in graphs
If $\mathsf{G}_{i}$ is a perfect graph for every $i\in[\ell-1]$ and $\mathsf{G}_{\ell}$ is the complement of the Schläfli graph, then we have:
  • $\Theta(\mathsf{G})=\alpha_{\mathrm{f}}(\mathsf{G})$ for every $i\in[\ell-1]$.
@@
Thus, the sufficient conditions of Theorem 3.3 hold, whereas the sufficient conditions of Theorem 3.2 do not hold.
@@ 4 The Shannon capacity of Tadpole graphs
$|\operatorname{\mathsf{V}}(\mathrm{T}({k},{\ell}))|=|\operatorname{\mathsf{E}}(\mathrm{T}({k},{\ell}))|=k+\ell.$
@@
The motivation to explore the Tadpole graphs comes from their similarity to the cycle graph, as graphs and as their equivalent DMCs, which is shown in Figure 3.
Figure 3: The DMC of Tadpole graphs (Left) and the DMC of cycles (Right)
@@
    $\vartheta(\mathrm{T}({k},{\ell}))=\left\lfloor\frac{k}{2}\right\rfloor+\left\lceil\frac{\ell}{2}\right\rceil.$
@@
    $\vartheta(\mathrm{T}({k},{\ell}))=\frac{k}{1+\sec\frac{\pi}{k}}+\frac{\ell}{2}.$
@@
First, if $k=3$ then the cycle $\operatorname{\mathsf{C}}_{3}$ is a clique of order $3$, and since the clique-cover number of $\operatorname{\mathsf{P}}_{\ell}$ is $\left\lceil\frac{\ell}{2}\right\rceil$ then
$\operatorname{\sigma}(\mathrm{T}({3},{\ell}))=1+\left\lceil\frac{\ell}{2}\right\rceil.$
By Lemma 4.1 and Theorem 2.21, a lower bound and an upper bound on $\vartheta(\mathrm{T}({k},{\ell}))$ coincide, which gives
$\vartheta(\mathrm{T}({3},{\ell}))=1+\left\lceil\frac{\ell}{2}\right\rceil.$
@@
$\operatorname{\sigma}(\mathrm{T}({k},{\ell}))=\left\lfloor\frac{k}{2}\right\rfloor+\left\lceil\frac{\ell}{2}\right\rceil,$
@@
$\vartheta(\mathrm{T}({k},{\ell}))=\left\lfloor\frac{k}{2}\right\rfloor+\left\lceil\frac{\ell}{2}\right\rceil.$
@@
$\leq\vartheta(\operatorname{\mathsf{C}}_{k}+\operatorname{\mathsf{P}}_{\ell}).$
Furthermore, by Theorems 2.23 and 2.24,
@@
$=\vartheta(\operatorname{\mathsf{C}}_{k}+\operatorname{\mathsf{P}}_{\ell}).$
@@
$\vartheta(\mathrm{T}({k},{\ell}))=\frac{k}{1+\sec\frac{\pi}{k}}+\frac{\ell}{2}.$
@@
$\Theta(\mathrm{T}({k},{\ell}))=\left\lfloor\frac{k}{2}\right\rfloor+\left\lceil\frac{\ell}{2}\right\rceil.$
@@
$\Theta(\operatorname{\mathsf{C}}_{k}+\operatorname{\mathsf{P}}_{\ell-1})\leq\Theta(\mathrm{T}({k},{\ell}))\leq\Theta(\operatorname{\mathsf{C}}_{k}+\operatorname{\mathsf{P}}_{\ell}).$
Furthermore, path graphs are universal, then, by Theorem 3.3,
$\Theta(\operatorname{\mathsf{C}}_{k}+\operatorname{\mathsf{P}}_{\ell-1})$ $=\Theta(\operatorname{\mathsf{C}}_{k})+\Theta(\operatorname{\mathsf{P}}_{\ell-1})$
$=\Theta(\operatorname{\mathsf{C}}_{k})+\frac{\ell}{2}$
$=\Theta(\operatorname{\mathsf{C}}_{k}+\operatorname{\mathsf{P}}_{\ell}).$
Hence,
$\Theta(\mathrm{T}({k},{\ell}))=\Theta(\operatorname{\mathsf{C}}_{k})+\frac{\ell}{2}.$
@@
$\leq\vartheta(\operatorname{\mathsf{C}}_{k})=\frac{k}{1+\sec\frac{\pi}{k}}.$
@@
$\leq\frac{k}{1+\sec\frac{\pi}{k}}+\frac{\ell}{2}.$
@@
$\sqrt{\alpha(\mathsf{G}_{1}\boxtimes\mathsf{G}_{1})}=\sqrt{26}=5.09902,$
$\sqrt[3]{\alpha(\mathsf{G}_{1}\boxtimes\mathsf{G}_{1}\boxtimes\mathsf{G}_{1})}=\sqrt[3]{136}=5.14256.$
Let $\mathsf{G}_{2}=\mathrm{T}({7},{6})$, then by Corollary 4.1 and the lower bound [36]
$\Theta(\operatorname{\mathsf{C}}_{7})\geq\sqrt[5]{\alpha(\operatorname{\mathsf{C}}_{7}^{5})}\geq\sqrt[5]{367},$
which yields
@@
The last example shows that the lower bound in Corollary 4.1 can be improved. The following result gives an improved lower bound.
@@
The upper bound was proved in Corollary 4.1. We next prove the improved lower bound. By Hales’ paper [28, Theorem 7.1], if $k\geq 3$ is odd, then
$\alpha(\operatorname{\mathsf{C}}_{k}^{2})=\left(\frac{k-1}{2}\right)^{2}+\left\lfloor\frac{k-1}{4}\right\rfloor,$
@@
$=\sqrt{\left(\frac{k-1}{2}\right)^{2}+\left\lfloor\frac{k-1}{4}\right\rfloor}.$
Thus, by combining the last inequality with Theorem 4.1, it follows that
@@
$\geq\sqrt{\left(\frac{k-1}{2}\right)^{2}+\left\lfloor\frac{k-1}{4}\right\rfloor}+\frac{\ell}{2}.$
The gap between the upper and lower bounds on $\Theta(\mathrm{T}({k},{\ell}))$ for $k\geq 7$ odd and $\ell\geq 0$ even satisfies,
@@
$<\frac{k+\ell}{2}-\frac{k+\ell-1}{2}=\frac{1}{2}.$ (69)
@@
Theorem 4.2 gives an improved lower bound that is based on a lower bound of the capacity of odd cycles that was constructed by Hales (see [28]). Since hales’ result, better lower bounds for odd cycles were constructed, and can be used to improve the lower bound in special cases. For better lower bounds on higher powers of odd cycles see [37, 38]. The next theorem shows an improvement to the bound in (68), for special cases using a result from [37].
@@
Let $n\geq 1$ and $d\geq 1$ be natural numbers and let $\ell\geq 5$ be an odd number. Define
$k=n2^{d}+2^{d-1}+1.$
@@
$\Theta(\mathrm{T}({k},{\ell}))=\frac{\ell}{2}+\Theta(\operatorname{\mathsf{C}}_{5}).$ (71)
And, by [37, Theorem 1.4],
$\Theta(\operatorname{\mathsf{C}}_{5})$ $\geq\sqrt[d]{\alpha(\operatorname{\mathsf{C}}_{k})}$
@@
In Theorem 4.3 a bound from [37, Theorem 1.4] was used. It is worth noting that in his paper, Bohman conjectured that the lower bound was the exact value of the independence number (see [37, Conjecture 1.5]). In a later paper of Bohman, with R. Holzman and V. Natarajan [38], a countably infinite subset of these values was confirmed, and by substituting $d=3$, which gives $k=8n+5$, it was proven that if $k=8n+5$ is a prime number, then the lower bound was the exact value of the independence number, i.e.
@@ 5 When the graph capacity is not attained by the independence number of any finite power?
In Section 4, we proved that if $k=3$, or if $k\geq 4$ is even, or if $k\geq 5$ and $\ell\geq 1$ are odd, then the capacity of the Tadpole graph $\mathrm{T}({k},{\ell})$ coincides with its independence number. Additional families of graphs also share this property; for example, by [11], the capacities of all Kneser graphs are equal to their independence numbers. It is, however, well known that not all graphs posses this property; e.g., $\Theta(\operatorname{\mathsf{C}}_{5})=\sqrt{\alpha(\operatorname{\mathsf{C}}_{5}^{2})}=\sqrt{5}$ and $\alpha(\operatorname{\mathsf{C}}_{5})=2$. The property that the Shannon capacity coincides with the square root of the independence number of the second (strong) power of the graph was proved, more generally for all self-complementary vertex transitive graphs [11] and all self-complementary strongly regular graphs [34], provided that the order of the graph is not a square of an integer. Families of graphs whose capacity is attained at a finite power that is strictly larger than $2$ are yet to be found. In this section, we explore sufficient conditions on a graph $\mathsf{G}$ that makes its Shannon capacity be unattainable by the independence number of any finite power of $\mathsf{G}$, i.e., we explore the conditions on $\mathsf{G}$ to have:
$\Theta(\mathsf{G})>\sqrt[k]{\alpha(\mathsf{G}^{k})}\quad\forall\,k\in\mathbb{N}.$ (74)
For example, the graph $\mathsf{G}=\operatorname{\mathsf{C}}_{5}+\mathsf{K}_{1}$ satisfies (74) (see Example 5.1).
@@
The following example provides a disconnected graph whose capacity is not attained at any finite power, the proof of this example uses an important concept that is later used in our first original approach. This example was shown by A. Wigderson and J. Zuiddam [22].
@@
$\Theta(\mathsf{C}_{5}\boxtimes\mathsf{K}_{1})=\Theta(\mathsf{C}_{5})=\Theta(\mathsf{C}_{5})\,\Theta(\mathsf{K}_{1}).$
@@
$\Theta(\mathsf{G})=\Theta(\mathsf{C}_{5})+\Theta(\mathsf{K}_{1})=\sqrt{5}+1,$
@@
$=c_{k}+d_{k}\sqrt{5}\notin\mathbb{N},$
@@
$c_{k}$ $=\sum_{\ell=0}^{\bigl\lfloor\frac{k}{2}\bigr\rfloor}\binom{k}{2\ell}\,5^{\ell}\in\mathbb{N}$
$d_{k}$ $=\sum_{\ell=0}^{\bigl\lfloor\frac{k-1}{2}\bigr\rfloor}\binom{k}{2\ell+1}\,5^{\ell}\in\mathbb{N}.$
It therefore follows that, for all $k\in\mathbb{N}$,
$\Theta(\mathsf{G})^{k}\neq\alpha(\mathsf{G}^{k})$
since the independence number is an integer. Consequently, (74) holds.
We next consider three approaches for the construction of families of graphs whose Shannon capacities satisfies the condition in (74). The first approach is due to Guo and Watanabe [10] with a simplified proof (the original proof from [10] is provided in Appendix C), and the second and third approaches of constructing such graph families are original.

### v2

@@ 3 The Shannon capacity of polynomials in graphs
Let $\mathsf{G}_{i}$ be a perfect graph for every $i\in[\ell-1]$, and let $\mathsf{G}_{\ell}$ be the complement of the Schläfli graph. Then,
  • $\Theta(\mathsf{G}_{i})=\alpha_{\mathrm{f}}(\mathsf{G}_{i})$ for every $i\in[\ell-1]$.
@@
Hence, the sufficient conditions of Theorem 3.3 hold, in contrast to those of Theorem 3.2.
@@ 4 The Shannon capacity of Tadpole graphs
$|\operatorname{\mathsf{V}}(\mathrm{T}({k},{\ell}))|=|\operatorname{\mathsf{E}}(\mathrm{T}({k},{\ell}))|=k+\ell.$ (96)
@@
The motivation for studying Tadpole graphs stems from their similarity to cycle graphs, both as graphs and in terms of their equivalent DMCs, as illustrated in Figure 3.
Figure 3: The DMCs of $\mathrm{T}({k},{\ell})$ (left plot) and $\operatorname{\mathsf{C}}_{k+\ell}$ (right plot).
@@
    $\vartheta(\mathrm{T}({k},{\ell}))=\left\lfloor\frac{k}{2}\right\rfloor+\left\lceil\frac{\ell}{2}\right\rceil.$ (98)
@@
    $\vartheta(\mathrm{T}({k},{\ell}))=\frac{k}{1+\sec\frac{\pi}{k}}+\frac{\ell}{2}.$ (99)
@@
First, if $k=3$, then $\operatorname{\mathsf{C}}_{3}$ is a clique of order $3$. Since the clique-cover number of $\operatorname{\mathsf{P}}_{\ell}$ is $\left\lceil\frac{\ell}{2}\right\rceil$, it follows that
$\operatorname{\sigma}(\mathrm{T}({3},{\ell}))=1+\left\lceil\frac{\ell}{2}\right\rceil.$ (100)
By Lemma 4.1 and Theorem 2.22, the lower and upper bounds on $\vartheta(\mathrm{T}({k},{\ell}))$ coincide, which gives
$\vartheta(\mathrm{T}({3},{\ell}))=1+\left\lceil\frac{\ell}{2}\right\rceil.$ (101)
@@
$\operatorname{\sigma}(\mathrm{T}({k},{\ell}))=\left\lfloor\frac{k}{2}\right\rfloor+\left\lceil\frac{\ell}{2}\right\rceil,$ (102)
@@
$\vartheta(\mathrm{T}({k},{\ell}))=\left\lfloor\frac{k}{2}\right\rfloor+\left\lceil\frac{\ell}{2}\right\rceil.$ (103)
@@
$\leq\vartheta(\operatorname{\mathsf{C}}_{k}+\operatorname{\mathsf{P}}_{\ell}).$ (104)
Furthermore, by Theorems 2.24 and 2.25,
@@
$=\vartheta(\operatorname{\mathsf{C}}_{k}+\operatorname{\mathsf{P}}_{\ell}).$ (105)
@@
$\vartheta(\mathrm{T}({k},{\ell}))=\frac{k}{1+\sec\frac{\pi}{k}}+\frac{\ell}{2}.$ (106)
@@
$\Theta(\operatorname{\mathsf{C}}_{k}+\operatorname{\mathsf{P}}_{\ell-1})\leq\Theta(\mathrm{T}({k},{\ell}))\leq\Theta(\operatorname{\mathsf{C}}_{k}+\operatorname{\mathsf{P}}_{\ell}).$ (110)
Furthermore, as path graphs are universal (see Theorem 2.24), it follows from Theorem 3.3 that
$\Theta(\operatorname{\mathsf{C}}_{k}+\operatorname{\mathsf{P}}_{\ell-1})$ $=\Theta(\operatorname{\mathsf{C}}_{k})+\Theta(\operatorname{\mathsf{P}}_{\ell-1})$ (111)
$=\Theta(\operatorname{\mathsf{C}}_{k})+\frac{\ell}{2}$ (112)
$=\Theta(\operatorname{\mathsf{C}}_{k}+\operatorname{\mathsf{P}}_{\ell}).$ (113)
Combining (110) and (113) gives equality (108). ∎
The next result follows from equality (108).
@@
$\leq\vartheta(\operatorname{\mathsf{C}}_{k})=\frac{k}{1+\sec\frac{\pi}{k}}.$ (117)
@@
$\leq\frac{k}{1+\sec\frac{\pi}{k}}+\frac{\ell}{2}.$ (118)
@@
$\sqrt{\alpha(\mathsf{G}_{1}\boxtimes\mathsf{G}_{1})}=\sqrt{26}=5.09902,$ (120)
$\sqrt[3]{\alpha(\mathsf{G}_{1}\boxtimes\mathsf{G}_{1}\boxtimes\mathsf{G}_{1})}=\sqrt[3]{136}=5.14256.$ (121)
Let $\mathsf{G}_{2}=\mathrm{T}({7},{6})$. By the lower bound in [26] on the Shannon capacity of the cycle $\operatorname{\mathsf{C}}_{7}$,
$\Theta(\operatorname{\mathsf{C}}_{7})\geq\sqrt[5]{\alpha(\operatorname{\mathsf{C}}_{7}^{5})}\geq\sqrt[5]{367},$ (122)
which implies by Corollary 4.1 and (122) that
@@
The last example shows that the lower bound in Corollary 4.1 is not tight. The following result provides an improved lower bound.
@@
The upper bound was proved in Corollary 4.1. We next prove the improved lower bound. By Hales’ result [32, Theorem 7.1], if $k\geq 3$ is odd, then
$\alpha(\operatorname{\mathsf{C}}_{k}^{2})=\left(\frac{k-1}{2}\right)^{2}+\left\lfloor\frac{k-1}{4}\right\rfloor,$ (125)
@@
$=\sqrt{\left(\frac{k-1}{2}\right)^{2}+\left\lfloor\frac{k-1}{4}\right\rfloor}.$ (126)
Thus, by combining (126) with Theorem 4.1, it follows that
@@
$\geq\sqrt{\left(\frac{k-1}{2}\right)^{2}+\left\lfloor\frac{k-1}{4}\right\rfloor}+\frac{\ell}{2}.$ (127)
The gap between the upper and lower bounds on $\Theta(\mathrm{T}({k},{\ell}))$ in (124), for an arbitrary odd $k\geq 7$ and even $\ell\geq 0$, satisfies
@@
$<\frac{k+\ell}{2}-\frac{k+\ell-1}{2}$ (128)
$=\frac{1}{2},$
where inequality (128) follows by dropping the term $\left\lfloor\frac{k-1}{4}\right\rfloor$ inside the square root of the first line, and since $\sec\frac{\pi}{k}>1$.
@@
Theorem 4.2 gives an improved lower bound that relies on a lower bound on the Shannon capacity of odd cycles that was constructed by Hales (see [32]). Since Hales’ result, better lower bounds for odd cycles were constructed, and can be used to improve the lower bound in special cases. For better lower bounds on higher powers of odd cycles see [27, 41]. The next theorem shows an improvement to the bound in (124), for special cases using a result from [27].
@@
Let $n,d\in\mathbb{N}$, let $\ell\geq 0$ be an even number, and define
$k=n2^{d}+2^{d-1}+1.$ (129)
@@
$\Theta(\mathrm{T}({k},{\ell}))=\frac{\ell}{2}+\Theta(\operatorname{\mathsf{C}}_{k}),$ (131)
and, by [27, Theorem 1.4],
$\Theta(\operatorname{\mathsf{C}}_{k})$ $\geq\sqrt[d]{\alpha(\operatorname{\mathsf{C}}_{k}^{d})}$
@@
The proof of Theorem 4.3 relies on the lower bound on $\alpha(\operatorname{\mathsf{C}}_{k}^{d})$ given in (4), as presented in [27, Theorem 1.4]. It is worth noting that in [27], Bohman conjectured that the lower bound is the exact value of the independence number (see [27, Conjecture 1.5]). In a subsequent paper by Bohman, Holzman and Natarajan [41], the conjecture was confirmed for a countably infinite subset of these values. In particular, by substituting $d=3$, which gives $k=8n+5$, it was proved in [41] that if $8n+5$ is a prime number, then the lower bound coincides with the exact value of the independence number. Specifically, by [41, Theorem 1], the independence number of the strong cubed power of the cycle graph $\operatorname{\mathsf{C}}_{8n+5}$ is given by
@@ 5 When the graph capacity is not attained by the independence number of any finite power?
In Section 4, we proved that if $k=3$, or if $k\geq 4$ is even, or if $k\geq 5$ and $\ell\geq 1$ are odd, then the capacity of the Tadpole graph $\mathrm{T}({k},{\ell})$ coincides with its independence number. Additional families of graphs also share this property; e.g., by [11], the capacities of all Kneser graphs are equal to their independence numbers. It is, however, well known that not all graphs posses this property; e.g., $\Theta(\operatorname{\mathsf{C}}_{5})=\sqrt{\alpha(\operatorname{\mathsf{C}}_{5}^{2})}=\sqrt{5}$ and $\alpha(\operatorname{\mathsf{C}}_{5})=2$. The property that the Shannon capacity coincides with the square root of the independence number of the second (strong) power of the graph was proved, more generally for all self-complementary vertex transitive graphs [11] and for all self-complementary strongly regular graphs [39], provided that the order of the graph is not a square of an integer. Families of graphs whose capacity is attained at a finite strong power that is strictly larger than $2$ are yet unknown. In this section, we explore sufficient conditions on a graph $\mathsf{G}$ such that its Shannon capacity is unattainable by the independence number of any finite strong power of $\mathsf{G}$, i.e.,
$\Theta(\mathsf{G})>\sqrt[k]{\alpha(\mathsf{G}^{k})},\quad\forall\,k\in\mathbb{N}.$ (134)
@@
The following example exhibits a disconnected graph whose capacity is not attained by the independence number of any strong finite power of the graph. The proof of this example relies on an important concept that is later employed in our first original approach. This example, appearing explicitly in [23], was essentially already contained in [10] and was later explicitly pointed out in the 1998 paper on zero-error information theory by Körner and Orlitsky [5] (at the end of Section IV), where it is attributed to Arikan.
@@
$\Theta(\mathsf{C}_{5}\boxtimes\mathsf{K}_{1})=\Theta(\mathsf{C}_{5})=\Theta(\mathsf{C}_{5})\,\Theta(\mathsf{K}_{1}).$ (135)
@@
$\Theta(\mathsf{G})=\Theta(\mathsf{C}_{5})+\Theta(\mathsf{K}_{1})=\sqrt{5}+1,$ (136)
@@
$=c_{k}+d_{k}\sqrt{5}\notin\mathbb{N},$ (137)
@@
$c_{k}=\sum_{\ell=0}^{\bigl\lfloor\frac{k}{2}\bigr\rfloor}\binom{k}{2\ell}\,5^{\ell}\in\mathbb{N},\qquad d_{k}=\sum_{\ell=0}^{\bigl\lfloor\frac{k-1}{2}\bigr\rfloor}\binom{k}{2\ell+1}\,5^{\ell}\in\mathbb{N}.$ (138)
It therefore follows that, for all $k\in\mathbb{N}$, $\Theta(\mathsf{G})^{k}\neq\alpha(\mathsf{G}^{k})$ since the independence number is an integer. Consequently, (134) holds.
The interested reader is referred to [42, Section 5], where Example 5.1 is studied in the context of zero-error variable-length coding.
We next consider three approaches for the construction of families of graphs whose Shannon capacities satisfies the condition in (134). The first approach is due to Guo and Watanabe [10] with a simplified proof (the original proof from [10] is provided in Appendix D), and the second and third approaches of constructing such graph families are original.